OpenStudy (czesc):

INSTANT FAN//MEDAL What is distance around the graph of r = 12 sin θ ?

OpenStudy (mathmale):

It appears that you simply deleted your previous post of this question, along with the info and questions I added. Why would you do that? Certainly this does not motivate me to spend more time with you, nor would it motivate anyone else to do so.

OpenStudy (czesc):

@mathmale I didn't delete it, I closed it, because I couldn't bump it again for 23:00 and I had no information regarding finding distance around a graph anywhere on the internet.

OpenStudy (samslays):

@czesc Don't let what @mathmale says get to you. He is supposed to be a moderator, but he is very rude

OpenStudy (anonymous):

I vote: rewirte in (x,y) coordinates, see it's a circle, and use the perimeter formula.

OpenStudy (anonymous):

the rewrite isn't to bad: multi by r to get ---> \[r^2=12r\sin(\theta)\] and recall that \[x^2+y^2=r^2 \text{ and } y=\sin(\theta)\]

OpenStudy (anonymous):

then complete the square to get standard form \[(x-x_0)+(y-y_0)=r^2\]

OpenStudy (mathmale):

Thanks for the compliment, @Samslays. @czesc: It may help you to define what you mean by "distance around the graph." This is decidedly non-standard mathematical language. Perhaps you mean "arc length?" or even "perimeter?"

OpenStudy (anonymous):

arc length is my bet.

OpenStudy (mathmale):

Also, I'd suggest you actually graph r= 12 sin theta. If, as another user seems to claim, this is a circle, then rewriting the given equation in rectangular coordinate form may make this problem seem quite easy.

OpenStudy (czesc):

@mathmale thank you for your response, I appreciate it as always, regarding distance around the graph, i'm just as confused as you are as I did not write the question.

OpenStudy (czesc):

Yes indeed it is a circle resting with it's bottom on the x-axis.

OpenStudy (anonymous):

Then arc length is easy! \[2\pi r\] if you want \[0<\theta\le2\pi\]

OpenStudy (czesc):

@Redcan @mathmale the options for response are this, but when I try to find arc length of polar curve r=12*sin(t) from t=0 to t=2pi i get 24 pi and the options are 6π 144π 36π 12π

OpenStudy (anonymous):

r= 6

OpenStudy (anonymous):

\[2*6*\pi\]

OpenStudy (anonymous):

\[x^2+(y-6)^2=6^2\] ----> circle of radius 6